A different manufacturing company began 2 new jobs during July. At the beginning of July, Job 227 was the only job in work in process inventory. There was no finished goods inventory. The cost of beginning work in process inventory for Job 227 consists of $1,500 in direct material cost, $2,000 in direct labor cost, and manufacturing overhead costs of $4,500. Total manufacturing overhead for the month was $16,054. The following information is available for costs added during the month of July:

job 227 job 228 job 229
Direct Material $935 $2850 $1300
Direct labor $1840 $3225 $1975
direct labor hours 184 310 204

Job 227 was completed and moved into finished goods inventory towards the beginning of the month. Later in the month, John 227 was sold. Job 228 was completed, but has yet to be sold. John 229 remains in production.

Required: Keep the following principles in mind as you follow the instructions below. First, costs for any job can only be in one place at the end of the month. Second, you will have to calculate the overhead rate for the month of July based upon the number of hours budgeted in production divided by the amount of manufacturing overhead budgeted. The annual budget for production was Budgeted Overhead of $180,000 and Budgeted Direct Labor Hours of 8,000. Third, the flow of funds is very important, but the detail on any job is only collected in the work in process inventory.

  1. Calculate the cost of the ending work in process inventory as of the end of the month.

  2. Calculate the cost of the finished goods inventory as of the end of the month.

  3. Calculate the cost of goods sold for the month.
Answer #1
Budgeted Production Overhead $      180,000
Budgeted Direct Labor Hours               8,000
Predetermined Production overhead rate/Direct Labor hour= $           22.50
Production Cost details for the Month of July Job 227 Job 228 Job 229 Total
Beginning WIP
Direct Materials $           1,500 $                   1,500
Direct Labor $           2,000 $                   2,000
Manufacturing overhead $           4,500 $                   4,500
A Total Beginning WIP $           8,000 $                   8,000
Costs Added during July
a Direct Materials $              935 $              2,850 $               1,300 $                   5,085
b Direct Labor $           1,840 $              3,225 $               1,975 $                   7,040
c Direct Labor Hrs                  184                      310                      204
d Manufacturing overhead @$22.5/Direct Labor Hr=c*22.5= $           4,140 $              6,975 $               4,590 $                 15,705
B Total Costs Added during July =a+b+d= $           6,915 $            13,050 $               7,865 $                 27,830
Total Costs of Jobs at Month end=A+B $         14,915 $            13,050 $               7,865 $                 35,830
Job status Sold in FG in WIP
Cost of Ending WIP at the end of July (Job 229)= $           7,865
Cost of FG Inventory at the end of July (Job 228)= $         13,050
Cost of Goods Sold during July (Job 227)= $         14,915